# Strong asymptotic giant branch stars’ spectral features in distant quiescent galaxies: Impact on galaxy evolution

### v1

@@ 4.5 Spectral fit
To estimate the stellar mass, age, and SFH of a galaxy, we adopt a custom IDL routine [68] to fit the final corrected 1D spectrum of each galaxy. The best-fitting results are derived by comparing the 1D spectrum with a range of composite stellar population templates by -minimisation. These templates were generated by combining a grid of different SSP models assuming a delayed exponentially declining SFH (, i.e., delayed- model). The characteristic timescale and age variations are the same in different models, shown in Table 4. To identify the redshift, the Calzetti attenuation law [69] was adopted, and the stellar metallicity was left free with lower and upper limits dependent on the model. To reduce the computational cost, we run the spectral fit twice. First, we run the spectral fit in a large range of redshift with a low-resolution grid (). After the most probable redshift peak was identified, we narrowed down the redshift range with a high-resolution grid () in the second run. The redshift boundary in Table 4 is an example of a parametric grid for D36123 in the second run. For object D36123 with the highest SNR, the spectral fit was repeated varying the wavelength range (see Table 1): (1) with the full rest-frame wavelength (0.3-2m, see Fig 10); (2) with the optical only (m); (3) with the NIR only (m); and (4) with the full rest-frame wavelength range but excluding 0.5–1m (see Fig. 11). For the two fainter galaxies, the spectral fit was only done in the full rest-frame wavelength. Since the medium resolution grating spectrum of object 9025 does not cover too well the age-sensitive spectral region around rest-frame 4000 break, we additionally considered the photometry in HST F606W and F814W bands to better constrain the overall fit.
@@
The best-fitting parameters with 1 errors[70] can be derived from the distribution of , which are listed in Table 2. Considering the large minimum reduced of object D36123, we derived the uncertainties (as listed in Table 1) scaling up all errors to obtain a 1.0. All stellar masses are reported for a Chabrier IMF [55], and are corrected for stellar losses and include remnants. Fig.10 shows the best-fit model for each galaxy using different models in red at the top of each panel, and the relative flux deviation at the bottom of each panel. The SNR-weighted averages of the relative deviations are labeled in the top-right corner of each bottom-panel. For object D36123 the models deviate from the spectrum at a few percent level (2.5% level for M13, and higher values for the other models). This is likely driven by systematics in the models (plus, possible residual calibration errors) which we can detect at the high SN ratio. Higher deviations correspond to observed features not matched by the models. For the other galaxies the relative flux deviations are larger and dominated by the noise in the data.
4.6 Additional details on models
@@
The C09 models allow several user options and have been modified through the years. The models used in this work were generated using the latest FSPS version 3.2 by adopting a Chabrier IMF, the BaSTI isochrones and the MILES stellar library, without considering dust extinction. We correct for the expected stellar losses and include remnants. The C09 models include TP-AGB spectra, extrapolated blueward with simpler linear slopes (as advocated by Lancon & Wood 2000 [13]) and redward with the Aringer et al. (2009) [71] synthetic carbon star spectral library. The spectra of the oxygen-rich TP-AGB stars are extrapolated with the latest version of the PHOENIX stellar spectral library (the BT-SETTL library). For the TP-AGB isochrones, there is no shift adopted in the () vs. (). For crude understanding of the effect of TP-AGB stars in the NIR mass-to-light ratio we show 1 Gyr SSP models at solar metallicity from the different libraries, in Fig.12.
| Parameters [units] | Min | Max | Step |
| 1.06 | 1.1 | 0.001 | |
| Age [Gyr] | 0.1 | 3.5 | 0.1 |
| 0.001, 0.004, 0.009, 0.05 | |||
| [Gyr] | 0.1 | 1.0 | 0.1 |
| 1.0 | 3.0 | 1.0 | |
| Av [mag] | 0 | 2.1 | 0.1 (Calzetti) |
| 0.3 | 2.5 | 0.1 (BC03) | |
| 0.3 | 2.0 | 0.1 (C09) | |
| 0.3 | 2.2 | 0.1 (M13) | |
| 0.5 | 2.0 | 0.1 (M05) | |

### v2

@@ 3.5 Spectral fit
To estimate the stellar mass, age, and SFH of a galaxy, we adopt a custom IDL routine[64] to fit the final corrected 1D spectrum of each galaxy. The best-fitting results are derived by comparing the 1D spectrum with a range of composite stellar population templates by -minimisation. These templates were generated by combining a grid of different SSP models assuming a delayed exponentially declining SFH (, i.e., delayed- model). The characteristic timescale and age variations (i.e., the time since the onset of star formation) are the same in different models, shown in Extended Data Table 1. To identify the redshift, the Calzetti attenuation law[65] was adopted, and the stellar metallicity was left free with lower and upper limits dependent on the model. To reduce the computational cost, we run the spectral fit twice. First, we run the spectral fit in a large range of redshift with a low-resolution grid (). After the most probable redshift peak was identified, we narrowed down the redshift range with a high-resolution grid () in the second run. The redshift boundary in Extended Data Table 1 is an example of a parametric grid for D36123 in the second run. For object D36123 with the highest SNR, the spectral fit was repeated varying the wavelength range (see Table 1): (1) with the full rest-frame wavelength (0.3-2m, see Extended Data Fig.4a-d); (2) with the optical only (m); (3) with the NIR only (m); and (4) with the full rest-frame wavelength range but excluding 0.5–1m (see Extended Data Fig.4e-h). For the two fainter galaxies, the spectral fit was only done in the full rest-frame wavelength. Since the medium resolution grating spectrum of object 9025 does not cover too well the age-sensitive spectral region around rest-frame 4000 break, we additionally considered the photometry in HST F606W and F814W bands to better constrain the overall fit.
@@
The best-fitting parameters with 1 errors[66] can be derived from the distribution of , which are listed in Table 2. Considering the large minimum reduced of object D36123, we derived the uncertainties (as listed in Table 1) scaling up all errors to obtain a 1.0. Stellar masses are reported for a Chabrier IMF [51] and are corrected for stellar losses and include remnants. Ages reported in Tables and discussed in the paper are mass-weighted. Extended Data Fig.4 shows the best-fit model for each galaxy using models in red at the top of each panel, and the relative flux deviation at the bottom of each panel. The SNR-weighted averages of the relative deviations are labelled at the top of each bottom panel. For object D36123 the models deviate from the spectrum at a few percent levels (2.5% level for M13, and higher values for the other models). This is likely driven by systematics in the models (plus, possible residual calibration errors) which we can detect at the high SNR. Higher deviations correspond to observed features not matched by the models. For the other galaxies, the relative flux deviations are larger and dominated by the noise in the data.
Fitting additional models to D36123: In addition to the models discussed in the main text, we also fitted three additional models to the galaxy D36123. This includes two fully empirical sets of models as in XSL[35] and E-MILES[36]. We also tested CB07 models (https://www.bruzual.org/cb07/), an update of BC03 that was meant to include strong TP-AGB (never formally published). Some descriptions of the recipes of these models are provided in a further section in these Methods. Their intrinsic dispersion and resolutions are also shown on the top panels of Extended Data Fig.3. Both XSL and E-MILES models have higher intrinsic resolutions than other models. The parameter grid settings in these three models are the same as those of the four main models, except for the metallicity due to available ranges (see Extended Data Table 1 for details). Their corresponding best-fit models are displayed in Extended Data Fig.5, accompanied by the best-fit redshift, mass-weighted age, stellar mass and reduced shown in Extended Data Table 2, only for the full-range fits. The XSL model does not cover the blueward of 3500Å, and includes two noisy regions corresponding to the low atmospheric transmission which are masked during the spectral fit (see grey-shaded regions of Extended Data Fig.5).
3.6 Selected features and indexes on D36123
We provide quantitative measurements of a range of features and indexes in D36123, chosen to be particularly representative as a test of the underlying contributing stellar phases, and also of models. To this aim, we have to take into account the fairly reduced spectral resolution of the PRISM spectrum of D36123, which triggers ad-hoc re-definition of some indexes even when similar indexes were already defined in the literature. However by measuring these indexes in the same way in matched-resolution best-fitting models as well in representative stellar phases (including C- and O-type TP-AGB[13, 14] and RGB[35]) also at matched resolutions, we can gain valuable scientific insights (see Fig.3).
We define two EW indexes, that are estimated with the equation:
where Fc is the continuum obtained by drawing a straight line (see the blue dashed lines in Fig.3a,b) between the two-side average continuum from to , estimated over 100Å (rest) regions (red crosses).
1) EW(0.5-0.53m) index: this includes MgI and MgII features, overlapping the MgII index, but also TiO and CN features. The continuum side windows are at the rest-frame 0.5-0.51m and 0.53-0.54m, while the index is integrated over rest 0.51-0.53m.
2) EW(0.9-1.0m) index: this covers the TiO+ZrO ‘valley’ and the CN0.92 edge. The continuum side windows are at the rest-frame 0.91-0.92m and 0.98-0.99m, while the index is integrated over rest 0.92-0.986m.
We further define two strength indexes as follows:
3) H-bump index. This broad bump, sensitive to the presence of cool stars, is the combination of the H- peak opacity at 1.6m and of H2O vapour absorptions around 1.4m and 1.9m. We adopt the literature definition[67]:
the Fcont pseudo-continuum flux (see the blue dashed line in Fig.3c) is estimated by connecting the midpoint of the two-side continuum windows at rest-frame 1.45-1.47m and 1.765-1.785m (red crosses). Fλ is integrated from 1.61 m to 1.67m.
4) CN1.1 index.
Among all CN edges, CN1.1 is the strongest detection in D36123. We define the CN1.1 index by the following formula:
where the pseudo-continuum has a normalization and slope that are both measured using a single window at rest wavelength bluer than the edge between 1.05-1.075m, 250Åwide. The need to define a slope is to allow distinguishing steeply declining continuum over a large range from genuine edge-like discontinuities. A positive (rising slope) is, however, reset to a flat continuum to avoid inflating artificially the index. We then extrapolate this pseudo-continuum over the range 1.1-1.12m where we integrate the spectrum (see Fig.3d).
3.7 Additional details on model input
@@
The BC03 models[8] are calculated with an isochrone synthesis algorithm, using Padova and Geneva tracks, plus extra ingredients to describe the TP-AGB phase and adopt both theoretical (BaSeL) and observational (STELIB and Pickles) libraries of stellar spectra, where the STELIB and Pickles libraries are extended to the shorter and longer wavelengths by using the BaSeL. The CB07 (and CB16) models (https://www.bruzual.org/) are updated versions of the BC03 model. The C09 models allow several user options and have been modified through the years. The models used in this work were generated using the latest FSPS version 3.2 by adopting a Chabrier IMF, the BaSTI isochrones and the MILES stellar library in the optical. The C09 models include TP-AGB spectra, extrapolated blueward with simpler linear slopes (as advocated by Lancon & Wood 2000[13]) and redward with the Aringer et al. (2009)[68] synthetic carbon star spectral library. The spectra of the oxygen-rich TP-AGB stars are extrapolated with the latest version of the PHOENIX stellar spectral library (the BT-SETTL library). For the TP-AGB isochrones, there is no shift adopted in the () vs. (). Maraston models (i.e., M05 and M13) utilise the fuel-consumption theorem for the treatment of post-main-sequence phases[3] and adopt empirical C-/O-rich spectra for the TP-AGB phase[14] and the BaSeL library for the other phases. The XSL model[35] utilises the PARSEC/COLIBRI isochrones including a TP-AGB contribution and adopting the X-shooter Spectral Library, which contains many evolved cool giants, covering the wavelength range from 3500 to 2.48m (shorter than other models). These cool giants with 4000 K are incorporated into the XSL model by using average spectra of static giants (i.e., RGB), O-/C-rich TP-AGB stars, binned by broad-band colour in the Vega system. The E-MILES model[36] adopts BaSTI stellar tracks for all phases, including the TP-AGB and the E-MILES library(http://miles.iac.es/pages/) for stellar spectra. All models assume a Chabrier IMF and are corrected to a solar metallicity in this work. For the crude understanding of the effect of TP-AGB stars in the NIR mass-to-light ratio, we show 1 Gyr SSP models around solar metallicity from the different libraries in Extended Data Fig.6. SSPs from CB16, Vazdekis, XSL, E-MILES, and CB07 are degraded to match the wavelength-dependent resolution of M13 for better visualization.
| Parameters [units] | Min | Max | Step | |
| 1.06 | 1.1 | 0.001 | ||
| 0.1 | 3.0 | 0.1 | ||
| [Gyr] | 3.5 | 5.0 | 0.5 | |
| 5.0 | 10.0 | 1.0 | ||
| 10.0 | 20.0 | 5.0 | ||
| 0.001, 0.004, 0.009, 0.05 | ||||
| [Gyr] | 0.1 | 1.0 | 0.1 | |
| 1.0 | 3.0 | 1.0 | ||
| Av [mag] | 0 | 2.1 | 0.1 (Calzetti) | |
| 0.3 | 2.5 | 0.1 (BC03) | ||
| Main | 0.3 | 2.0 | 0.1 (C09) | |
| 0.3 | 2.2 | 0.1 (M13) | ||
| 0.5 | 2.0 | 0.1 (M05) | ||
| 0.3 | 1.6 | 0.1 (XSL) | ||
| Additional | 0.3 | 2.5 | 0.1 (E-MILES) | |
| 0.3 | 2.5 | 0.1 (CB07) | ||
† indicates the time since the onset of star formation. The solar metallicity 0.02 is adopted.
| Rest-frame | Property | XSL | E-MILES | CB07 |
| redshift | 1.084 | 1.083 | 1.080 | |
| ] | 1.866 | 2.065 | 1.119 | |
| age [Gyr] | 1.556 | 1.934 | 0.991 | |
| full | 1.319 | 0.506 | 0.975 | |
| (0.3–2m) | Av | 0.052 | 0.000 | 0.975 |
| [Gyr] | 0.159 | 0.385 | 0.007 | |
| SFR | 0.073 | 0.460 | 0.000 | |
| SFRbest/SFRpeak | 0.001 | 0.090 | 0 | |
| 55.5 | 46.1 | 117.0 |
| H-bump index | CN1.1 index | EW(Mg+TiO+C2) | EW(CN0.92+TiO+ZrO) | ||
| Spec | Bin | 1.61-1.67m | 1.1m | 0.5-0.53m | 0.9-1m |
| D36123 | – | -0.0840.002 | -0.1160.003 | 9.6390.486 | 33.5690.944 |
| BC03 | – | -0.048 | -0.0140.027 | 5.246 | 6.934 |
| C09 | – | -0.143 | -0.0360.019 | 4.933 | 12.716 |
| M13 | – | -0.115 | -0.0510.009 | 3.908 | 11.311 |
| M05 | – | -0.147 | -0.0770.004 | 5.841 | 11.009 |
| XSL | – | -0.158 | 0.0070.001 | 6.793 | 19.967 |
| E-MILES | – | -0.081 | -0.0110.004 | 3.369 | 16.151 |
| CB07 | – | -0.028 | -0.0340.008 | 3.462 | 8.453 |
| (I-K)=2.01 | -0.124 | -0.014 | 20.466 | 12.771 | |
| RGBa | (I-K)=2.54 | -0.144 | -0.043 | 22.407 | 22.139 |
| (I-K)=3.41 | -0.143 | -0.031 | 23.838 | 30.149 | |
| (I-K)=2.11 | -0.304 | -0.030 | 22.849 | 28.061 | |
| O-rich TP-AGBa | (I-K)=3.19 | -0.378 | -0.112 | 23.624 | 47.078 |
| (I-K)=4.88 | -0.690 | -0.105 | – | – | |
| bin1 | -0.281 | -0.061 | – | – | |
| O-rich TP-AGBb | bin5 | -0.370 | -0.01 | – | – |
| (I-K)=2.61 | -0.218 | -0.549 | 43.837 | 228.035 | |
| C-rich TP-AGBa | (I-K)=3.65 | -0.408 | -0.498 | 72.194 | 257.859 |
| (I-K)=4.46 | -0.491 | -0.358 | – | – | |
| bin1 | -0.338 | -0.385 | – | – | |
| C-rich TP-AGBb | bin3 | -0.299 | -0.309 | – | – |
The superscript a represents the average spectra of cool giants binned by broad-band colour (in Vega system) from the XSL simple stellar population models[35], while superscript b are binned O- and C-rich spectra, observed by Lançon & Wood[13], and Lançcon & Mouhcine[14], respectively. Each feature of star spectra is measured after smoothing and rebinning to match the PRISM resolution around the feature position.
@@
The Supplementary Table 1 lists all proposed NIR features in the object D36123.
| This work | Literature | ||||
| Type | No. | [Å] | specific name | [Å] | index name |
| 1 | 5395-5555 | CN0.55 | – | – | |
| 2 | 5900-6080 | CN0.6 | – | – | |
| 3 | 6921-7060 | CN0.7 | – | – | |
| edge | 4 | 7740-7890 | CN0.78 | – | – |
| (C-rich) | 5 | 9110-9290 | CN0.92 | – | – |
| 6 | 10830-11040 | CN1.1 | 10780-11120[69, 18, 40] | CN11 | |
| 7 | 13755-13885 | CN1.38 | – | – | |
| 8 | 17660-17745 | CN1.77 | – | – | |
| 9 | 5045-5285 | MgI+MgII+TiO(0)+C2? | 5069-5197[67, 40] | Mg1/Mg2/Mgb | |
| 10 | 5560-5730 | TiO(0)+C2 | – | – | |
| 11 | 5900-5990 | TiO(1) | 5937-5994[67, 40] | TiO1 | |
| 12 | 6170-6355 | TiO(0)+C2 | 6191-6272[67, 40] | TiO2 | |
| absorption | 13 | 7114-7310 | TiO(0) | – | – |
| (O-rich) | 14 | 7790-7985 | VO B-X (0) | – | – |
| 15 | 8455-8645 | TiO(0)+VO B-X(-1)+CaII | 8476-8700[40] | CaT1/CaT2/CaT3 | |
| 16 | 8830-9015 | TiO(0) | – | – | |
| 17 | 9465-9655 | TiO(-1)+ZrO? | 9200-9500[18, 40] | CN/TiO/ZrO | |
| 18 | 10458-10675 | VO A-X (0) | 10470-10650[18, 40] | VO | |
NIR features for which we propose identification in the rest-frame spectrum of object D36123, include (C-rich) spectral edges and (O-rich) absorptions. These rest-frame wavelength range of specific features are approximate owing to the fairly low resolution of the spectrum, and also because the broadness of the features make it harder to clearly define a center. Some specific features defined by previous studies[67, 69, 18, 40] are listed in the right two columns for comparison.